A 160 MW solar PV power plant produces 248 GWh/yr. Determine the equivalent hours of operation per year at rated power.
Answer #1

Energy produced by the plant in one second = 160 MW

The energy produced by the plant in one hour = 160 * 3600 MW

Energy production per =248 *103 * 3600 MW /year

so the number of hours /year at rated power = 248 *103 * 3600/ 160 * 3600 = 1550 hrs in one year

ans 1550 hours